OpenStudy (anonymous):

The sum of the squares of two consecutive positive even integers is 340. Find the integers.

OpenStudy (freckles):

what number comes right after n?

OpenStudy (anonymous):

That's the full question. That's all this test gave me. It says "The sum of the squares of two consecutive poistive even integers is 340. Find the Integers"

OpenStudy (anonymous):

Positive**

OpenStudy (freckles):

yes i know but what comes after n?

OpenStudy (anonymous):

I don't know!

OpenStudy (freckles):

like for instance if i asked you: what comes after 1, you would say 1+1

OpenStudy (freckles):

if i said what comes after 2, you would say 2+1

OpenStudy (freckles):

if i said what comes after 3, you would say 3+1

OpenStudy (anonymous):

So would it be n+n? or n+1??

OpenStudy (freckles):

I'm adding 1 to each of those not n

OpenStudy (freckles):

so n+1

OpenStudy (freckles):

n,n+1 these are consecutive positive numbers

OpenStudy (freckles):

of course we are assuming n is positive

OpenStudy (freckles):

now it says the sum of the squares of these numbers

OpenStudy (anonymous):

those aren't numbers.. Those are letter o_O

OpenStudy (freckles):

so that means we will adding these numbers squares

OpenStudy (anonymous):

Do you know the answer to this or what??

OpenStudy (freckles):

\[n^2+(n+1)^2=\] It says the sum of these squares =340 so what should come after that = sign?

OpenStudy (freckles):

I'm not gonna give you the answer

OpenStudy (freckles):

I will guide you

OpenStudy (freckles):

You need to solve this for n \[n^2+(n+1)^2=340 \]

OpenStudy (freckles):

Do you know how to solve this?

OpenStudy (anonymous):

I got \[\chi \approx 12.5288 , -13.5288\]

OpenStudy (anonymous):

Which is in no way the answer -__-

OpenStudy (anonymous):

Your problem does not have integer solution.

OpenStudy (freckles):

@eliassaab is right

OpenStudy (anonymous):

What do you mean?

OpenStudy (freckles):

that positive number 12.5288 is not an integer

OpenStudy (anonymous):

Well duhh. That's why i need help??

OpenStudy (freckles):

wait your question said even?

OpenStudy (anonymous):

I got the answer. its' 12 and 14. Thanks for the "help"

OpenStudy (freckles):

So if 2 is the first even integer then what is the next even integer? 2,2+2 right? if 4 is the first even integer then what is the next even integer? 4,4+2 ----- we add 2 ----- if n is the first even integer, then what is the next even integer?

OpenStudy (freckles):

why do you have to say it like that @DeaneNicole ? I was just trying to help

OpenStudy (anonymous):

No no no no no no no no no i didn't mean it like that!! Sorry /:

OpenStudy (freckles):

Ok ok. Its all cool.

OpenStudy (anonymous):

Alright.

OpenStudy (freckles):

I'm sorry I miss the word even

OpenStudy (freckles):

all those tiny words have meaning lol

OpenStudy (anonymous):

You have to solve (2n)^2 + (2n+2)^2 =340 n =6 so 12 and 14

OpenStudy (anonymous):

@eliassaab we already got it.

OpenStudy (freckles):

and @eliassaab this would have worked too \[n^2+(n+2)^2=340 \]

OpenStudy (anonymous):

\[ 12^2+14^2=144+196=340 \]

OpenStudy (anonymous):

Sorry, I did not see your post above @DeaneNicole